OpenStudy (otherworldly):

Solve for x: 4 over 5 x + 4 over 3 = 2x

OpenStudy (3mar):

Do you mean\[\frac{ 4 }{ 5x }+\frac{ 4 }{ 3 }=2x\]?

OpenStudy (3mar):

or \[\frac{ 4 }{ 5 }x+\frac{ 4 }{ 3 }=2x\]?

OpenStudy (otherworldly):

that one yup

OpenStudy (3mar):

Which "that"?

OpenStudy (otherworldly):

the second one XD sorry

OpenStudy (otherworldly):

Can you help me please?

OpenStudy (otherworldly):

@3mar ?

alones (alones):

Alright subtract \(\ 2x\) from both sides

alones (alones):

Can you do that?

alones (alones):

\[\frac{ 4 }{ 5 }x+\frac{ 4 }{ 3 }-2x =2x−2x\]

alones (alones):

\[\frac{ -6 }{ 5 }x+\frac{ 4 }{ 3 }=0\]

OpenStudy (otherworldly):

is there another way because that way doesn't seem correct...?

alones (alones):

Now try subtracting \(\frac{ 4 }{ 3 }\) from each side

alones (alones):

waitt but we're not done yet.

OpenStudy (otherworldly):

ok

alones (alones):

Anyways when we subtract 4 by 3..we need to get rid of that \[\frac{ -6 }{ 5 }x+\frac{ 4 }{ 3 }-\frac{ 4 }{ 3 }\] Which would equal \[0-\frac{ 4 }{ 3 }\]

OpenStudy (otherworldly):

-4/3

alones (alones):

Not that's not the answer.

OpenStudy (otherworldly):

so -6/5x = -4/3

OpenStudy (otherworldly):

ik I'm just following along

alones (alones):

So ^^yes that At last multiply both sides by \[\frac{ 5 }{ (-6) }\]

jhonyy9 (jhonyy9):

@AloneS sorry but i think that there above you need multiply the 2x by 5 not by 3

OpenStudy (otherworldly):

@AloneS shouldn't i multiply both sides by -6/5?????

jhonyy9 (jhonyy9):

just bc. 4/5 is the coeficient of x and not 4/3

alones (alones):

\[(\frac{ 5 }{ -6 }) *(\frac{ -6 }{ 5}x)=(\frac{ 5 }{ -6 })*\frac{ -4 }{ 3 }\]

alones (alones):

Yes mybad lol

OpenStudy (otherworldly):

?

OpenStudy (otherworldly):

This is confusing me

OpenStudy (otherworldly):

Can someone explain to me a different way please?

jhonyy9 (jhonyy9):

4 4 ---x + --- = 2x subtract from both sides 4/3 so what will get ? 5 3

alones (alones):

Anyways after dividing that you would get \(\frac{ 10 }{ 9 }\)

OpenStudy (otherworldly):

4/5x = 2x - 4/3

jhonyy9 (jhonyy9):

4 4 ---x = 2x - --- and now subtract from both sides 2x and will get 5 3 4 ---x -2x = - 4/3 5 (4x-10x)/5 = (-4)/3 cross multiply and will get 3(-6x) = -20 so do you can continue it now ?

OpenStudy (otherworldly):

ok thank you @AloneS i will follow along @jhonyy9 because he is doing it a different way and i need to follow along because i feel like it would be easier for me

OpenStudy (otherworldly):

how did u get (4x-10x)/5

jhonyy9 (jhonyy9):

on the left side there are 4 ---x -2x yes ? so make common denominator and will get (4x-10x)/5 yes ? 5

alones (alones):

Lol sorreh @OtherWorldly for making it difficult LOL Alright. (:

OpenStudy (otherworldly):

y r u putting a minus sign when it is supposed to be " / "

jhonyy9 (jhonyy9):

no where ?

OpenStudy (otherworldly):

(4x-10x)/5

jhonyy9 (jhonyy9):

there is (4x-10x) divide by 5

jhonyy9 (jhonyy9):

bc when you make common denominator so multiply the 2x by 5 and will get 10x/5 but there are 4 10x ---x - ----- = (4x-10x)/5 yes ? 5 5

OpenStudy (otherworldly):

yeah srry i must of gotten lost in the middle

jhonyy9 (jhonyy9):

how or why this is easy

jhonyy9 (jhonyy9):

so on the left side will get -6x/5

OpenStudy (otherworldly):

ok

jhonyy9 (jhonyy9):

and on the right side what are there ?

OpenStudy (otherworldly):

-20

OpenStudy (otherworldly):

?

OpenStudy (otherworldly):

right ??

jhonyy9 (jhonyy9):

no -4/3

jhonyy9 (jhonyy9):

so we have than -6x -4 ---- = ---- 5 3

OpenStudy (otherworldly):

so multiply 5 on both sides right

jhonyy9 (jhonyy9):

not just this cross multiply and what will get ?

OpenStudy (otherworldly):

-18x= -20 ???

jhonyy9 (jhonyy9):

exactly and now ?

OpenStudy (otherworldly):

9/10

jhonyy9 (jhonyy9):

there are -2*9x = -2*10 so what can you make on the every two sides ?

OpenStudy (otherworldly):

-18x=-20.....

OpenStudy (otherworldly):

x= 9/10

jhonyy9 (jhonyy9):

what you see there is a common coeficient ?

jhonyy9 (jhonyy9):

-2 so divied both sides by -2 and will get ?

jhonyy9 (jhonyy9):

9x=10 right ?

OpenStudy (otherworldly):

yes

jhonyy9 (jhonyy9):

so from this how you get the x ?

jhonyy9 (jhonyy9):

courage

OpenStudy (otherworldly):

x= 1.1 ?

jhonyy9 (jhonyy9):

how you ve got this ?

OpenStudy (otherworldly):

omg i don't know i divided 10 by 9

OpenStudy (otherworldly):

i made it into decimal but in fraction form its 10/9 its the same thing

jhonyy9 (jhonyy9):

np this is easy way to solve it 9x = 10 divide both sides by 9 and will get x = 10/9

OpenStudy (otherworldly):

ok thank you soooooo much i took up too much of ur time but i appreciate it

jhonyy9 (jhonyy9):

anytime just tag me was my pleasure hope you understood it right now good luck bye bye

OpenStudy (otherworldly):

i hope so too thx bye